x-intercept of a(3,5) and b(-2,-3)
Answer by stanbon(75887)

x-intercept of a(3,5) and b(-2,-3)
----
slope = (5--3)/(3--2) = 8/5
-----
y-int = ?
y = mx + b
5 = (8/5)3 +b
(25/5) = (24/5) + b
b = 1/5
============
Equation:
y = (8/5)x + (1/5)
----
x-intercept = ?
Let y = 0
The (8/5x) = -(1/5)
x = 1/8
=================
Cheers,
Stan H.
=================